# Testing Low-Resource Language Support in LLMs Using Language Proficiency Exams: the Case of Luxembourgish

### v2

@@ 2.5 Prompt Design
We experimented with prompts of different lengths and degrees of complexity to find a prompt that yields the best for solving the language exams. See Appendix C for the full prompt. Regarding the prompts for the headline generation and description summary tasks, we reuse the prompts used by Plum et al. (2025) for every model we investigate. The prompts used for evaluating the LuxGen tasks that uses the metric requirements described in Section 2.4 can be found in Appendix D.
@@ 3.1 RQ1: How well do LLMs perform in Luxembourgish language tests?
Table 2 shows the overall results of the LLMs we tested. See Appendix E for a breakdown of results across the four categories. For a more convenient way of exploring the results, we also release an interactive web app with the results 555which can be found at https://applux-nnjdvksvbpgovzux7dtqts.streamlit.app/.
We observe a clear trend where the large models typically perform well above average. Claude 3.5 Sonnet achieves the highest scores across every test except for the C2 test for which DeepSeek-R1:671B performs best. Among the small models, many models show promise to be adequate for further fine-tuning on Luxembourgish data despite their small size as they either approach or slightly surpass the 50% mark on the easier tests. Gemma2:9B performs best on the easier tests, but we deem GLM4 to be the best small model overall as it gets close to Gemma2’s scores for the easier tests and tends to perform best on the harder tests.
There is also a trend appearing of small models outperforming medium-sized models, with the latter performing worse than guessing at random. Llama3.1:70B is the overall winner among the medium-sized models, however, it performs significantly worse than GLM4, and indeed most of the small models.
Furthermore, as expected, performances drop for higher CEFR levels as the tests become increasingly more difficult, with one notable exception. Curiously, most models achieve a higher performance on the B2 test than on the B1 test. Specifically, they perform surprisingly well on the Reading Comprehension questions, however, it is unclear as to why those questions appear to be easier for LLMs to solve.
@@ 3.2 RQ2: In which cases do LLMs fail to understand Luxembourgish?
In order to answer this question, we analyse trends and patterns relating to the four categories mentioned in Section 2.2.1. Table 3 contains the performances of the studied language models per category. From the results, we note that overall, LLMs perform almost consistently best in vocabulary questions while grammar appears to be the hardest category. Furthermore, although the Reading Comprehension and Conversation Comprehension questions are nominally different, there is no statistical difference between the performances of the LLMs in either category (according to a paired t-test). However, when taking into account the size categories of the LLMs, large models perform significantly better in the reading comprehension category than conversation comprehension questions.
In addition to the coarse-grained ones, we created fine-grained categories for the questions in the language tests. Table 4 shows the categories as well as the number of questions that fit into the categories. We present the results of the best-performing models in Table 5. Due to space issues, we put the entire table in Appendix F. The categorisation was done in a supervised fashion using Claude 3.7 Sonnet and a separate manually annotated language exam. We provide the descriptions for each category as well as the prompt for annotation in Appendices G and H. We manually reviewed the automatically annotated questions to ensure correctness.
model Size Vocab. Grammar RC CC
@@
Qwen 2 0.5 26.5 21.8 19.4 21.9
Llama 3.2 1 10.3 4.8 1.3 1.3
Qwen 2 1.5 34.8 32.1 23.9 25.2
StableLM 2 1.6 14.8 13.9 5.2 4.5
Gemma 2 2 29.0 24.8 27.7 29.7
Llama 3.2 3 33.5 27.9 38.7 32.3
Phi 3 3.8 27.7 26.1 34.8 34.8
Phi 3.5 3.8 14.8 4.8 11.6 8.4
Qwen 2 7 34.2 32.7 45.2 38.7
Llama 2 7 12.3 6.7 14.8 26.5
WizardLM 2 7 20.6 26.7 29.0 28.4
Mistral 7 30.3 27.3 36.8 38.1
Aya-23 8 30.3 28.5 39.4 35.5
Llama 3.1 8 32.9 34.5 50.3 39.4
Llama 3 8 34.2 36.4 38.7 38.1
DeepSeek-R1 8 38.1 30.9 32.9 37.4
GLM 4 9 50.3 24.8 51.6 48.4
Gemma 2 9 43.9 24.8 41.3 34.2
Mistral-Nemo 12 37.4 27.9 40.0 35.5
StableLM 2 12 27.1 14.5 31.0 27.7
Llama 2 13 18.1 14.5 12.9 14.8
Phi 3 14 36.1 37.6 38.7 37.4
Medium-sized Models ($>15B\&\leq 200B$)
Mistral-Small 22 7.7 1.2 0.0 0.6
Gemma 2 27 12.3 8.5 2.6 2.6
Aya-23 35 9.0 1.8 1.3 3.2
Command-R 35 5.2 0.6 0.6 1.9
Alfred 40 1.3 1.8 0.0 1.3
Mixtral 8x7 0.6 0.6 0.0 0.0
Llama 2 70 0.0 1.2 0.0 0.6
Llama 3 70 22.6 12.7 1.9 3.2
Llama 3.1 70 26.5 14.5 2.6 5.2
DeepSeek-R1 70 20.6 18.2 0.6 2.6
Qwen 2 72 12.3 7.3 2.6 5.2
Command-R+ 104 14.8 3.0 2.6 2.6
Mistral-Large 123 21.3 6.7 1.3 1.3
Mixtral 8x22 1.9 1.8 0.0 0.6
WizardLM 2 8x22 7.7 4.8 0.6 1.3
Large Models ($>200B$)
Llama 3.1 405 64.5 43.6 62.6 53.5
DeepSeek-R1 671 93.5 73.9 84.5 76.1
ChatGPT 3.5 unk. 92.9 72.1 87.1 80.0
Claude 3.5 Sonnet unk. 89.0 81.8 88.4 81.9
ChatGPT 4o unk. 89.7 71.5 84.5 80.6
ChatGPT 4o mini unk. 78.1 55.2 72.9 65.8
Gemini 2.0 Flash unk. 83.2 67.3 83.2 79.4
LeChat unk. 61.3 34.5 40.0 45.8
Category name Number
V_idiom 21
V_noun 84
V_verb 44
V_adjective 23
V_native 6
G_gender 37
G_pronouns 20
G_verbs 55
G_adjective 9
G_tense 50
G_past 21
G_present 13
G_subjunctive 14
G_modal 7
G_preposition 54
G_cases 39
RC_TF 152
RC_Simple 72
RC_Complex 86
Table 4: Statistics for fine-grained categories for exam questions.
It is to note that for this categorisation, we do not differentiate between reading comprehension and conversation comprehension, and thusly group them together.
Op ___ Weiere kann een de Wanter Schlittschong fueren
with the possible answers being:
zougefruer, zougefruerene, zougefruerenen, zougefréieren
For the reading comprehension questions, we split them into three categories: true/false, simple, and complex. Simple questions do not require an understanding of the entire given text excerpt, and the correct answer for those questions can be found either explicitly written in the text or by understanding a small portion of the text. True/false and complex questions require a model to fully understand the given text excerpt, the main difference being that True/False questions are binary while complex ones have three or four possible answers. As such, the latter questions are considered harder. This assumption is indeed confirmed by the results as the large models seem to perform best for simple questions.
It is to note, that these observations largely hold true for the best-performing small models, but performances tend to fluctuate for worse-performing models as well as for medium-sized models.
@@
3.3 RQ3: Are there correlations between performing well in a language exam and performing well in a language generation task?
@@
Description - Grammar 0.664 0.765 0.782 0.385
Description - Adequacy 0.645 0.724 0.622 0.267
Headline - Grammar 0.486 0.552 0.093 0.352
Headline - Adequacy 0.644 0.598 0.269 0.284
Table 6: Pearson Correlation Coefficient between the overall exam scores and LuxGen task performances
For this research question, we first average out the performances of every LLM on all six language exams, resulting in their overall exam performance score. We then let the LLMs solve both of the LuxGen tasks, and evaluate the outputs using an LLM-as-a-judge. Our LLM of choice is Claude 3.7 Sonnet Anthropic (2025). After manually verifying its capabilities to solve the tasks at hand, we deemed it an adequate LLM to evaluate the experiments as it solved both tasks well and with minimal spelling mistakes. For both tasks and both metrics we employ (see Section 2.4), we calculate mean performance scores for each LLM. Figure 2 shows the performances of all LLMs on the LuxGen tasks ($y$-axis) plotted against their respective overall exam performances ($x$-axis). See Appendix I for detailed results. We observe the formation of well-defined clusters, each corresponding to a different LLM size category which remains relatively consistent across each task and metric. Generally, the LuxGen performances of the small and large LLMs appear to strongly positively correlate with their exam performances. However, the medium-sized LLMs seem to not follow the same trend. Despite their generally low exam scores, many of them perform better on LuxGen tasks than other LLMs that reached high exam scores. However, when considering each cluster individually, we largely observe strong positive correlations that indeed show higher LuxGen performances for higher exam scores.
To confirm our observations, we calculate Pearson’s Correlation Coefficients (PCCs) for each set of performances (see Table 6). Across the entire set of LLMs, the PCCs are positive and weak to moderate, while within clusters, the PCCs typically show that performances are highly positively correlated. One notable exception is the large category of LLMs for the News Headline generation task, where there appears to be either a weak positive correlation (for adequacy) or no correlation at all (for grammar and spelling).
(a) Description Task - Grammar and Spelling
(b) Description Task - Adequacy
(c) Headlines Task - Grammar and Spelling
(d) Headline Task - Adequacy
Figure 2: Performance on Language Exams vs Performance on the LuxGen Tasks
@@ 4 Discussion
Generally, we conclude that language exams can indeed be a useful tool to assess the generation capabilities of LLMs. For the NLP tasks we tested, we can to a degree predict if an LLM performs well given how well it scored on the exams. However, there are some caveats to keep in mind:
First, for a considerable number of LLMs, these exams are not an adequate method. LLMs that have between 15 billion and 200 billion parameters scored consistently low on the exams, yet they performed well in the LuxGen tasks. Secondly, it is important to consider for what type of task an LLM is considered. Language exams are used to evaluate the general proficiency in a language across multiple disciplines. As such, the overall exam score can be dragged down by irrelevant categories of questions, e.g., an LLM used to detect hate speech in an online forum does not necessarily need to have a perfect grasp on obscure grammar rules. In such a case, it is preferable to adjust the exam to only include certain question categories. Thirdly, for this study, we did not do any additional fine-tuning on Luxembourgish data which might have altered the performances of the tested LLMs to different degrees. It is possible that certain LLMs adapt better when fine-tuned on a Luxembourgish instruction tuning dataset.
5 Related Works
5.1 Testing LLMs on Exams
5.2 Testing LLMs on Language Exams
However, compared to related studies, we do not only assess LLM performance on language exams alone, but we also explore their usefulness to predict their performance in general NLP tasks.
@@ 6 Conclusion
In this study, we investigated the performance of a large number of recent LLMs on official language exams, and assessed whether or not they serve as a useful tool to determine the capabilities of language models. We found that many models struggle with even the easier exams, with only the largest models scoring well in every exam. We also delved into dividing the exam questions into fine-grained categories to assess which types of questions are harder to solve for LLMs, finding that questions requiring language-exclusive knowledge as well as questions with similar answers pose problems to many models. Finally, we determined that using such exams can indeed be used as an indicator for text generation and understanding capabilities of LLMs, with some caveats to consider.
@@ 7 Acknowledgments
A Limitations
While we are confident about the findings and conclusions of our studies, there are still certain limitations to consider. First, while we aimed to include a comprehensive list of current LLMs in our experiments, it is a nearly impossible task considering the frequent release of new models. We even had to scrap parts of our experiments that were done using Gemini 1.5 Flash as it became inaccessible after the release of the 2.0 model. We are also fully aware that many new model versions have been released that we did not consider for this study. We however remain confident that our selection of LLMs gives a good overview of available models. Another limitation relates to the relatively small number of generation tasks we considered for RQ3. As they are essentially both summarisation tasks, it is possible that our conclusions do not apply to other types of NLP tasks. This could be a useful research direction for future work. Finally, it is possible that our conclusions are not generalisable to other languages since we are limiting ourselves to Luxembourgish. Again, studying if we can apply our methodology to other low-resource languages would be a worthwhile research direction that can benefit other language communities.

### v4

@@ 3.5 Prompt Design
We experimented with prompts of different lengths and degrees of complexity to find a prompt that yields the best result for solving the language exams. See Appendix C for the full prompt. Regarding the prompts for the Headline Generation and Short Description summary tasks, we reuse the prompts used by Plum et al. (2025) for every model we investigate. The prompts used for evaluating the LuxGen tasks that use the metric requirements described in Section 3.4 can be found in Appendix D.
3.6 Computational Setup and Hyperparameters
LLMs are either accessed externally via API key or stored locally on our HPC infrastructure where we have access to three NVIDIA Tesla V100 32GB GPUs. We use Ollama  [^4] https://github.com/ollama/ollama to prompt the LLMs.
For all our experiments, models and metrics, we use the default hyperparameters. The exceptions are the models solving the language exams, for which we set the temperature to 0 to reduce randomness of the answers.
@@ 4.1 RQ1: How well do LLMs perform in Luxembourgish language tests?
Table 2 shows the overall results of the LLMs we tested as well as a baseline determined as the average performance for a model that randomly guesses the answers of the exams. See Appendix E for a breakdown of results across the four categories. For a more convenient way of exploring the results, we also release an interactive web app with the results.
We observe a clear trend where the large models typically perform well above average. Claude 3.5 Sonnet achieves the highest scores across every test except for the C2 test for which DeepSeek-R1:671B performs best. Among the small models, many models show promise to be adequate for further fine-tuning on Luxembourgish data despite their small size as they perform better than the random baseline and either approach or slightly surpass the 50% mark on the easier tests. We deem Gemma3:12B the best small model as it performs best in all exams out of the models in the small category.
Medium-sized models generally perform consistently better than small models. Gemma3:27B is the overall winner among the medium-sized models, performing significantly better than all of its counterparts except for the C2 exams where it ties with the QWQ and Mistral-Large models.
Furthermore, as expected, performances drop for higher CEFR levels as the tests become increasingly more difficult, with one notable exception. Most models achieve a higher performance on the B2 test than on the B1 test. Specifically, they perform surprisingly well on the Reading Comprehension questions, warranting a closer examination of the composition of the exams.
@@ 4.2 RQ2: In which cases do LLMs fail to understand Luxembourgish?
model Size Vocab. Grammar RC $C$
@@
Qwen 2 0.5 25.7 19.7 21.2 22.6
Llama 3.2 1 21.3 23.7 20.5 18.8
Gemma 3 1 27.4 37.8 41.7 39.9
Qwen 2 1.5 29.6 28.3 25.0 25.2
StableLM 2 1.6 23.2 28.5 5.8 13.5
EuroLLM 1.7 24.0 20.0 16.7 18.1
Gemma 2 2 34.4 29.7 29.5 34.8
Llama 3.2 3 31.7 28.3 32.1 35.5
Phi 3 3.8 27.8 25.3 29.5 38.0
Phi 3.5 3.8 17.3 20.0 15.4 19.3
Gemma 3 4 46.2 32.9 43.6 47.1
Qwen 2 7 35.8 31.2 35.9 47.1
Llama 2 7 30.7 26.7 35.3 34.8
WizardLM 2 7 35.9 32.3 28.8 32.8
Mistral 7 31.2 28.3 34.6 38.1
Aya-23 8 31.7 29.5 32.1 37.4
Llama 3.1 8 30.1 33.5 39.7 44.4
Llama 3 8 34.5 33.7 35.3 38.6
DeepSeek-R1 8 40.6 34.5 40.4 41.3
GLM 4 9 50.3 27.3 43.6 52.8
EuroLLM 9 40.0 32.0 35.9 35.5
Gemma 2 9 57.3 34.1 44.9 43.8
Mistral-Nemo 12 36.2 27.7 34.6 36.7
StableLM 2 12 35.1 17.1 30.8 31.7
Gemma 3 12 63.1 32.4 54.5 59.9
Llama 2 13 25.6 26.6 21.2 21.3
Phi 3 14 34.0 35.8 32.7 38.6
Phi 4 14 46.2 36.7 45.5 45.8
Medium-sized Models ($15B\&\leq 200B$)
Mistral-Small 22 41.3 29.5 34.6 50.3
Gemma 2 27 60.1 36.7 57.1 62.0
Gemma 3 27 71.1 51.7 60.9 64.5
QWQ 32 59.5 38.1 41.7 56.0
Aya-23 35 39.6 30.7 40.4 40.0
Command-R 35 39.4 34.4 44.2 41.3
Alfred 40 34.3 32.3 34.6 36.8
Mixtral 8x7 47.1 27.3 36.5 40.7
Llama 2 70 28.8 28.3 36.5 43.1
Llama 3 70 53.0 35.6 40.4 43.8
Llama 3.1 70 60.4 46.0 45.5 50.9
DeepSeek-R1 70 64.5 42.2 44.9 60.6
Qwen 2 72 51.6 36.5 48.7 51.5
Command-R+ 104 46.5 31.1 40.4 42.5
Mistral-Large 123 64.9 41.1 42.3 52.2
Mixtral 8x22 49.4 40.5 42.3 50.9
WizardLM 2 8x22 40.4 37.7 41.7 49.0
Large Models (>200B>200B)
Llama 3.1 405 63.4 43.0 54.5 59.3
DeepSeek-R1 671 92.2 75.2 74.4 79.9
ChatGPT 3.5 unk. 91.6 72.1 73.7 85.8
Claude 3.5 Sonnet unk. 87.8 82.4 76.9 86.4
ChatGPT 4o unk. 88.3 71.5 73.1 84.5
ChatGPT 4o mini unk. 76.7 55.4 62.2 71.5
Gemini 2.0 Flash unk. 82.0 68.1 73.7 83.8
LeChat unk. 59.7 34.7 32.1 48.4
Table 3: Coarse-grained category results of language exams. Bold numbers indicate highest performance for each size category.
In order to answer this question, we analyse trends and patterns relating to the four categories mentioned in Section 1. Table 3 contains the performances of the studied language models per category. From the results, we note that overall, LLMs perform almost consistently best in vocabulary questions while grammar appears to be the hardest category. Furthermore, although the Reading Comprehension and Conversation Comprehension questions are nominally different, there is no statistical difference between the performances of the LLMs in either category (according to a paired t-test). However, when taking into account the size categories of the LLMs, large models perform significantly better in the conversation comprehension category than reading comprehension questions which might be due to the tested LLMs being specifically instruction-tuned to hold conversations.
In addition to the coarse-grained ones, we created fine-grained categories for the questions in the language tests (see Appendix G for more information on the categories). We present the results of the best-performing models in Table 4. Due to space issues, we put the entire table in Appendix F.
The categorisation was done in two steps: first, we did an automatic categorisation in a supervised fashion using Claude 3.7 Sonnet and category explanations as well as a separate manually annotated language exam that were added to the prompt. We provide the descriptions for each category as well as the prompt for annotation in Appendices G and H. In a second step, we manually reviewed the automatically annotated questions to ensure correctness. It is to note that for this categorisation, we do not differentiate between reading comprehension and conversation comprehension, and thusly group them together.
@@
Op ___ Weiere kann een de Wanter Schlittschong fueren
with the possible answers being:
zougefruer, zougefruerene, zougefruerenen, zougefréieren
For the reading comprehension questions, we split them into three categories: true/false, simple, and complex. Simple questions do not require an understanding of the entire given text excerpt, and the correct answer for those questions can be found either explicitly written in the text or by understanding a portion of the text. True/false and complex questions require models to fully understand the given text excerpt, the main difference being that True/False questions are binary while complex ones have three or four possible answers. As such, the latter questions are considered harder. This assumption is indeed confirmed by the results as the large models seem to perform best for simple questions.
It is to note, that these observations largely hold true for small and medium-sized models except for simple reading comprehension questions which was often the hardest category for those models.
4.3 RQ3: Are there correlations between performing well in a language exam and performing well in an NLG task?
@@
Claude-G 0.651 0.735 0.291 0.837
Claude-A 0.648 0.585 0.301 0.794
METEOR 0.427 0.450 0.089 0.667
BERTScore 0.407 0.116 0.266 0.622
(a) Pearson Correlation Coefficients
small medium large all
Claude-G 6.127 7.743 2.171 10.938
Claude-A 6.072 5.145 2.255 9.327
METEOR 3.369 3.600 0.639 6.390
BERTScore 3.179 0.837 1.972 5.676
(b) T-Scores
small medium large all
Claude-G 1.299E-07 3.667E-10 0.0346 5.552E-15
Claude-A 1.583E-07 4.318E-06 0.0285 1.312E-12
METEOR 0.0014 0.0007 0.5259 4.998E-08
BERTScore 0.0025 0.4064 0.0541 6.578E-07
(c) P-Values
Table 5: Correlation tests between the overall exam scores and Headline Generation task performances.
(a) Claude Score - Grammar and Spelling
(b) Claude Score - Adequacy
(c) BERTScore
(d) METEOR Score
Figure 2: Performance on language exams vs performance on the Headline Generation task.
For this research question, we first average out the performances of every LLM on all six language exams, resulting in their overall exam performance score. We then let the LLMs solve both of the LuxGen tasks, and evaluate the outputs using an LLM-as-a-judge (LaaJ). Our LLM of choice is Claude 3.7 Sonnet Anthropic (2025). After manually verifying its capabilities to solve the tasks at hand, we deemed it an adequate LLM to evaluate the experiments as it solved both tasks well and with minimal spelling mistakes. In addition to the LaaJ-metrics, we employ the METEOR and BERTScore metrics to measure the similarity between LLM outputs and the corresponding ground truths which also allows us to assess the usefulness of our novel metrics (see Appendix J). For both tasks and all metrics we employ, we calculate mean performance scores for each LLM. Figure 2 shows the performances of all LLMs on the Headline Generation task ($y$-axis) plotted against their respective overall exam performances ($x$-axis). See Appendix K for plots for the Short Description task and Appendix I for detailed results for both tasks. Generally, the LuxGen performances of the LLMs appear to strongly positively correlate with their exam performances. This is true for all metrics, although the spread and average performance is lower for the non-LaaJ metrics.
To confirm our observations, we calculate Pearson’s Correlation Coefficients (PCCs) for each set of performances (see Table 5). For most cases, we note moderate to strong PCCs. Some notable exceptions are the medium and large categories of LLMs with BERTScore and METEOR metrics, respectively, where there appears to be either a weak positive correlation or no correlation at all. We further calculate corresponding t-scores and p-values, and show that our findings are significant in most cases for $\alpha=0.05$ except for the aforementioned exceptions and BERTScore for the large category. We repeat the same significance tests for the Short Description task (see Appendix K) and find consistently high positive PCCs for all metrics, significant at $\alpha=0.05$.
@@ 5 Discussion
Generally, we conclude that language exams can indeed be a useful tool to assess the generation capabilities of LLMs. For the NLP tasks we tested, we can to a certain degree predict if an LLM performs well given how well it scored on the exams. However, there are some caveats to keep in mind:
Firstly, it is important to consider for what type of task an LLM is considered. Language exams are used to evaluate the general proficiency in a language across multiple disciplines. As such, the overall exam score can be dragged down by irrelevant categories of questions, e.g., an LLM used to detect hate speech in an online forum does not necessarily need to have a perfect grasp on obscure grammar rules. In such a case, it is preferable to adjust the exam to only include certain question categories. Secondly, for this study, as there was no instruction tuning data available for Luxembourgish to the best of our knowledge, we did not do any additional fine-tuning which might have altered the performances of the tested LLMs to different degrees. It is possible that certain LLMs adapt better when fine-tuned on an appropriate instruction tuning dataset first.
@@ 6 Conclusion
In this study, we investigated the performance of a large number of recent LLMs on official language exams, and assessed whether or not they serve as a useful tool to determine the capabilities of language models. We found that many models struggle with even the easier exams, with only the largest models scoring well in every exam. We also delved into dividing the exam questions into fine-grained categories to assess which types of questions are harder to solve for LLMs, finding that questions requiring language-exclusive knowledge as well as questions with similar answers pose problems to many models. Finally, we determined that using such exams can indeed be used as an indicator for language generation and understanding capabilities of LLMs, with some caveats to consider.
@@ 7 Acknowledgments
8 Limitations
While we are confident about the findings and conclusions of our studies, there are still certain limitations to consider. First, while we aimed to include a comprehensive list of current LLMs in our experiments, it is a nearly impossible task considering the frequent release of new models. We even had to scrap parts of our experiments that were done using Gemini 1.5 Flash as it became inaccessible after the release of the 2.0 model. We are also fully aware that many new model versions have been released that we did not consider for this study. We however remain confident that our selection of LLMs gives a good overview of available models. A limitation that might have affected the performance on the proficiency exams is the availability of such exams in the training data. While the exams we received from INLL are not publicly available and we do not believe they could have been retrieved from a webcrawl or similar means, LLMs could have seen exams from other languages or simpler language skill placement tests in Luxembourgish. Another limitation relates to the relatively small number of generation tasks we considered for RQ3. As they are essentially both summarisation tasks, it is possible that our conclusions do not apply to other types of NLP tasks. This could be a useful research direction for future work. In addition, we only used a small sample of those two summarisation tasks due to the large number of LLMs we considered. Furthermore, as the data that was used to create those tasks are publicly available, it is possible that at least a subset of the studied LLMs saw that data during training which would have influenced their performance in those tasks. Finally, it is possible that our conclusions are not generalisable to other languages since we are limiting ourselves to Luxembourgish. Again, studying if we can apply our methodology to other low-resource languages would be a worthwhile research direction that can benefit other language communities.
A Number and Distribution of Open Questions in Language Exams
Category A1 A2 B1 B2 C1 C2 Total
Vocabulary 0 0 13 13 14 12 52
Grammar 0 0 0 17 14 25 56
Total 0 0 13 30 28 37 108
Table 6: Number of open questions in language exams